### v2

# ToMiE: Towards Explicit Exoskeleton for the Reconstruction of Complicated 3D Human Avatars

Abstract
In this paper, we highlight a critical yet often overlooked factor in most 3D human tasks, namely modeling complicated 3D human with with hand-held objects or loose-fitting clothing. It is known that the parameterized formulation of SMPL is able to fit human skin; while hand-held objects and loose-fitting clothing, are difficult to get modeled within the unified framework, since their movements are usually decoupled with the human body. To enhance the capability of SMPL skeleton in response to this situation, we propose a growth strategy that enables the joint tree of the skeleton to expand adaptively. Specifically, our method, called ToMiE, consists of parent joints localization and external joints optimization. For parent joints localization, we employ a gradient-based approach guided by both LBS blending weights and motion kernels. Once the external joints are obtained, we proceed to optimize their transformations in $SE(3)$ across different frames, enabling rendering and explicit animation. ToMiE manages to outperform other methods across various cases with hand-held objects and loose-fitting clothing, not only in rendering quality but also by offering free animation of grown joints, thereby enhancing the expressive ability of SMPL skeleton for a broader range of applications. The code is available at https://github.com/Yifever20002/ToMiE.
Figure 1: We show two complex scenarios in 3D human modeling: with hand-held objects and loose-fitting clothing, which cannot be accurately represented by the standard SMPL model. Our ToMiE can realize adaptive growth to enhance the representation capability of SMPL without needing for time-consuming case-specific customization, achieving state-of-the-art results in both rendering and human (including complex scenarios) animation.
1 Introduction
3D human reconstruction endeavors to model high-fidelity digital avatars based on real-world characters for virtual rendering and animating, which has been of long-term research value in areas such as gaming, virtual reality (VR), and beyond. Traditional methods, such as SMPL [36], achieve human body parameterization by performing principal component analysis (PCA) on large sets of 3D scanned meshes, allowing for the fitting of a specified identity. Recent neural rendering techniques have enabled implicit digital human modeling guided by Linear Blend Skinning (LBS) and SMPL skeleton, realizing lifelike rendering and animating from video inputs.
The neural-based 3D human rendering has been empowered by the cutting-edge technique, 3D Gaussian Splatting (3DGS) [25], for its real-time and high-quality novel view synthesis performance. By representing 3D gaussians under the canonical T-pose and utilizing the pre-extracted SMPL skeleton in the observation space, 3D human rendering results can be obtained from novel views in any frame. This stream of approaches proves high quality in rendering 3D humans that conform to the SMPL paradigm (e.g., avatars in tight-fitting clothing). However, we raise concerns regarding its ability to handle complicated human modeling involving hand-held objects or loose-fitting clothing.
In Fig. 1, we show two cases to illustrate the limitations of current 3D human modeling. On the one hand, the movements of hand-held objects, e.g., feather duster, are highly decoupled from the human body and thus cannot be represented by the SMPL. On the other hand, characters shot in-the-wild are dressed in clothing with high complexity in dynamics, rather than the tight-fitting clothing configured under strict experimental conditions. These complicated scenarios break the existing methods’ paradigm by assuming that surface of avatars should be bounded to the motion warping of SMPL skeleton in the same way as the human skin is. Rooted in the aforementioned scenarios, SMPL fails to accurately fit such 3D human models.
To this end, we break through the limitations of modeling complicated 3D human gaussians by maintaining an extended joint tree from SMPL skeleton. Although existing SMPL model has the potential capability to manually customize additional skeleton information, time-consuming and case-by-case adjustments are necessary in this case. To overcome this issue, we extend the SMPL skeleton with additional joints for each individual case adaptively. The growth is performed in an explicit and adaptive manner, and enables the fitting of complicated 3D human avatars, offering high-quality rendering with hand-held objects and loose-fitting clothing (e.g., novel view synthesis results in Fig. 1).
The main challenge of extending the SMPL skeleton is to determine where and how to grow additional joints. To avoid the unnecessary memory consumption and potential overfitting caused by an arbitrary growth, we first determine which of the joints are supposed to serve as parent joints by a localization strategy. We have empirically observed that parent joints requiring growth witness larger backpropagation gradients in their associated gaussians due to underfitting. However, determining the association of gaussians with different joints is non-trivial, as the SMPL’s LBS blending weights may not work with the human with hand-held objects or clothing. To more precisely define such association, we introduce the concept of Motion Kernels based on rigid body priors and combine them with LBS weights, resulting in more accurate gradient-based localization. After growth, we adaptively maintain an extended joint tree and update the extra joints by optimizing two MLP decoders for joints positions and rotations. The proposed method, termed as ToMiE, allows for explicit rendering and animation of complicated human avatars represented by extended joints.
By experiments on complicated cases of the DNA-Rendering dataset [7], ToMiE exhibits state-of-the-art rendering quality while maintaining the animatability that is significant for downstream productions. To summarize, our contributions are three-fold as follows:
  • 1)
    ToMiE, a method for creating an enhanced SMPL joint tree via an adaptive growth strategy, which is able to decouple complicated parts from the human body, thereby achieving state-of-the-art results in both rendering and explicit animation on target cases;
  • 2)
    a hybrid assignment strategy for gaussians utilizing LBS weights and Motion Kernels, combined with gradient-driven parent joints localization, to guide the growth of external joints;
  • 3)
    a joints optimization approach fitting local rotations across different frames while sharing joints positions.
2 Related Work
2.1 SMPL-based Human Mesh Avatars
Most of the recent success in digital human modeling can be attributed to the contributions of the SMPL [36] series, which parameterizes the human body as individual shape components and motion-related human poses through 3D mesh scanning and PCA. The pose blend shapes in SMPL describe human body deformations as a linear weighted blending of different joint poses, significantly improving the efficiency of editing and animating digital humans. Furthermore, it has been widely adopted for human body animation, thanks to the methods [9, 49] of estimating SMPL parameters from 2D inputs. Despite their wide range of applications, SMPL and its family still suffer from inherent limitations. Since the originally scanned 3D meshes are skin-tight, according to which the pose blend shapes are learned, the model is unable to handle significantly outlying meshes, such as human with hand-held objects and clothing like skirts. Although this could potentially be solved by auto-rigging [20, 42, 55, 1, 28, 54], it requires 2D/3D shapes as inputs, which is not accessible in human videos.
2.2 Neural Representation for 3D Human
Methods based on neural representations, such as NeRF [38] and 3DGS [25], have also been playing an important part in digital human reconstruction for their high-quality rendering capabilities. Early NeRF-based methods [52, 45, 27, 6, 13, 3, 4, 5, 10, 11, 12] aim to reconstruct human avatars by inputting monocular or multi-view synchronized videos. [50] enforce smooth priors based on neural Signed Distance Function (SDF) to obtain more accurate human geometry. Recent breakthroughs [32, 35, 60, 30, 17, 46, 29, 18, 40, 24, 16, 31, 34, 57, 26, 39, 21, 58] rely on 3DGS, enabling faster and more accurate rendering. All these methods register the T-pose in a canonical space and use LBS weights to guide the rigid transformations.
2.3 Rendering & Editing of Intricate 3D Human
We revisit the methods attempting to implicitly improve modeling of complicated 3D human. Animatable NeRF [44] defines a per-frame latent code to capture appearance variations across each frame. Simulation methods [48, 2] physically construct simple clothing but struggle with complex one and objects. [29, 14] additionally register global latent bones to compensate for the limitations in clothing rendering. They fail to explicitly decouple the clothing from human body, making precise control infeasible. Another stream [6, 18] leverages a human poses sequence as contexts to resolve appearance ambiguities. But correspondingly, they require a sequence of human poses for animation, adding to the challenges of editing. Moreover, these methods struggle to fit the object-level pose independent of the human poses sequence, such as hand-held items. We also note that some works [15, 37] introduce diffusion-based generative methods to enhance the realism of garment rendering, but these methods are restricted by the traditional SMPL and overlook the editing of complex clothing. Although SMPLicit [8] enables implicit interpolation of clothing types, it remains dependent on SMPL’s LBS process to generate the observed mesh. This limitation prevents localized explicit animation of loose-fitting clothing and restricts its application to external objects. Rendering-based human reconstruction method [19, 33, 53, 23, 22, 57, 59] can not achieve animation. Our approach enables explicit decoupling of hand-held objects and clothing with the human body by extending the SMPL joint tree, allowing for high-quality rendering and explicit animation in complicated scenarios.
3 Preliminaries
3.1 SMPL(-X) Revisited
Pre-trained on scanned meshes, the SMPL(-X) family [36, 43] employs a parameterized model to fit human bodies of different shapes and under different poses. The human mesh in each frame evolves from a canonical human mesh and is controlled by shape and pose parameters. Specifically, a 3D point $\bm{x}_{c}$ on the canonical mesh will be warped to obtain point in the observation space as
$\bm{x}_{o}=\sum_{k=1}^{K}\omega_{k}(\bm{x}_{c})(R_{k}(\bm{r}^{0})\bm{x}_{c}+t_{k}(\bm{j}^{0},\beta)),$ (1)
where $K$ is the total number of joints, $R_{k}$ is per-joint global rotation controlled by local joint rotations $\bm{r}^{0}$, and $t_{k}$ is per-joint translation controlled by joint positions $\bm{j}^{0}$ and human shape $\beta$. Notice that linear blending weight $\omega_{k}$ is a function of $\bm{x}_{c}$, which is regressed from large human assets.
The main issue with this LBS-based prior model is that it can only model tight-fitting avatars conforming to the SMPL(-X) paradigm, making it unsuitable for modeling complex human clothing in more generic cases. What even more challenging is its inability to handle hand-held objects that are fully decoupled from the human pose.
3.2 Human Gaussians Revisited
Human gaussians achieve high-quality real-time human rendering by combining the SMPL prior with 3DGS as the representation. The SMPL model naturally obtains the T-pose (i.e., all human poses are identity transformations) mesh in the canonical space and then mesh vertices are used to initialize the canonical gaussian units. Each gaussian is defined as
$G(\bm{x})=\frac{1}{(2\pi)^{\frac{3}{2}}|\bm{\Sigma}|^{\frac{1}{2}}}e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})},$ (2)
where $\bm{\mu}$ is the 3D gaussian center, and $\bm{\Sigma}$ is the 3D covariance matrix, which will be further decomposed into learnable rotation $\bm{R}$ and scale $\bm{S}$. Now we have $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$, which is performed by optimizing a quaternion $\bm{r}_{g}$ for rotation and a 3D vector $\bm{s}_{g}$ for scaling . Each gaussian is further assigned with color $c$ and opacity $\alpha$.
Once the 3D gaussians in the canonical space are obtained, each position $\bm{\mu}$ will be warped to the observation space according to Eq. 1. Next, the 3D gaussians of each frame are projected into 2D gaussians, followed by tile-based rasterization. Color of each pixel can be calculated by blending $N$ ordered gaussians following
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (3)
The 3D gaussians can be optimized and updated by adaptive density control, which primarily includes cloning, splitting, and pruning. Cloning and splitting are guided by gradients to control the number of gaussians, while pruning removes empty gaussians based on current opacity $\alpha$. The supervision upon human gaussians is derived from multi-view or monocular videos, enabling high-quality rendering and avatar animation.
4 Methods
Figure 2: The pipeline of ToMiE. We initialize the gaussians in the canonical space with a standard SMPL vertices.  (Sec. 4.4) We apply Linear Blend Skinning (LBS) to the gaussian position and utilize a network for rotation and scale correction. During the warmup phase, Adaptive LBS only utilizes the original SMPL skeleton. After adaptive growth, it further includes the newly grown external skeleton. Gaussian rasterization and gradients backpropagation.  (Sec. 4.1, Sec. 4.2) We employ a gradient-based parent joints localization method, and a motion kernel to optimize the gradient assignment process.  (Sec. 4.3) We maintain an extra joint book with MLPs, which generates explicit human pose, enabling the decoupling and explicit animating of hand-held objects and loose-fitting clothing.
Fig. 2 illustrates ToMiE’s adaptive joint growth and gaussian training strategy. Our goal is to extend the SMPL skeleton to handle complicated human scenarios. However, the abuse of growth can lead to unnecessary computational and memory overheads. For an efficient adaptive growth, we first propose a localization strategy of parent joints to ensure that only necessary joints are grown. Furthermore, we explicitly define the hand-held joints in the $SE(3)$ space and optimize them end-to-end through an MLP, ensuring alignment with the original skeleton of SMPL. The extended skeleton can thus support rendering and explicit animation. To address the limitation of LBS in guiding gaussian attributes apart from the positions, we further fine-tune the rotation and scale during training with a deformation field to achieve better non-rigid warping. Next, we will elaborate on these modules and the training strategy.
4.1 Motion Kernels-guided Joint Gradient Accumulation
A quantitative metric needs to be identified to determine whether a joint requires growth. We notice that, due to the poor fitting ability of existing SMPL-based human gaussian, it will leave larger gradients in complex human regions (e.g., the hand-held object in Fig. 2 ). In other words, identifying joints with larger accumulated gradients can help to indicate which parent joints are more likely to require growth. Bounded with each gaussian, the accumulated gradients will be first assigned to their corresponding joints. Let $g=\|(g_{x},g_{y},g_{z})\|_{2}$ represent the L2 norm of the gradient at each gaussian position. The gradients accumulation $g_{J}$ for the $k$-th joint can then be computed according to
$g_{J_{k}}=\frac{\sum_{i\in N}\omega_{k}(\bm{x}_{c})g}{\sum_{i\in N}\omega_{k}(\bm{x}_{c})}.$ (4)
We use $\bm{x}_{c}$ to stand for gaussian position in canonical space and $\bm{x}_{o}$ in the observation space.
It is worth emphasizing that $\omega_{k}$ here is a weight term determining the assignment of a gaussian to the $k$-th joint. Under the paradigm of SMPL representation, this weight corresponds to the LBS weight $\omega_{\text{lbs0}}$, and guides the rigid transformation of vertices on the SMPL mesh according to the human pose. To account for the differences between the human mesh with clothing and the vanilla SMPL mesh, [17] calculate $\omega_{\text{lbs}}$ by keeping using the LBS weight prior $\omega_{\text{lbs0}}$ and adding an extra learnable network $\Phi_{\text{lbs}}$ for fine-tuning. This formulation (with index $k$ omitted) can be summarized as
$\omega_{\text{lbs}}(\bm{x}_{c})={\omega_{\text{lbs0}}}(\text{NN}_{1}(\bm{x}_{c},\bm{V}))+\Phi_{\text{lbs}}(\bm{x}_{c}),$ (5)
where $\text{NN}_{1}$ stands for top-1 nearest-neighbor search algorithm and $\bm{V}$ represents the canonical standard SMPL vertices. It will be, however, clarified by us, that this nearest neighbor-based assigning method will no longer be feasible in cases with handheld objects and complex clothing.
In Fig. 3, we present an example of misclassification by the nearest-neighbor search algorithm in Eq. 5. Due to the lack of human topology constraints in the canonical space, incorrect classification of hand-held objects can occur, as shown in Fig. 3 (a). This part, following the naive nearest-neighbor search, would be assigned to the leg by mistake. Fig. 3 (b) shows the points belonging to the hand joint (the correct parent joint where the handheld object should grow). Yet being solely guided by LBS weights results in the voids by misclassification.
Figure 3: Principle of the Motion Kernel. Relying solely on LBS weights will bring the misclassification in the canonical space shown in (a) to the observation space in (b), resulting in voids. Our proposed motion kernel focuses on motion-dependent priors in the observation space, offering better robustness and being less sensitive to misclassifications in the canonical space. This aids for point assignment process in parent joints localization.
To mitigate this issue, we propose a more robust assignment method based on motion priors, which we call Motion Kernels. Specifically, in the observation space, the motion kernel of each point $\bm{x}_{o}$ with respect to the $k$-th joint position $\bm{j}_{k}$ is defined based on the changes in their pairwise Euclidean distances through all input $N$ frames, following
$\text{MK}(\bm{x}_{c},\bm{j}_{k})=\frac{1}{N}\sum_{i=1}^{N}\left(\left\|{\bm{x}_{o}}_{i}-{\bm{j}_{k}}_{i}\right\|_{2}-\mu\right)^{2},$ (6)
and
$\mu=\frac{1}{N}\sum_{i=1}^{N}\left\|{\bm{x}_{o}}_{i}-{\bm{j}_{k}}_{i}\right\|_{2}.$ (7)
Our motion kernel (MK) reflects the relative motion between each gaussian and joint. A smaller MK indicates that the pair of gaussian and joint is relatively stationary to each other, signifying a stronger association, while a larger MK suggests a higher degree of relative motion, indicating a weaker association. We further represent the assignment weight reflected by the MK as $\omega_{\text{MK}}(\bm{x}_{c})=\text{Normalize}_{k}(\text{MK}^{-1}(\bm{x}_{c},\bm{j}_{k}))$, and the final assignment weight in Eq. 4 (with index k omitted) becomes
$\omega(\bm{x}_{c})=\lambda\omega_{\text{MK}}(\bm{x}_{c})+(1-\lambda)\omega_{\text{lbs}}(\bm{x}_{c}),$ (8)
where $\lambda$ is a hyperparameter to balance the MK weight and LBS weight. Note that we do not completely abandon the LBS weight, because the MK cannot differentiate the association of points on either side of a joint (“misdirect” in Fig. 3), requiring the LBS weight to compensate.
4.2 Gradient-based Parent Joint Localization
By combining Eq. 4 and Eq. 8, the parent joints $\bm{J}_{s}\subseteq\bm{J}$ that require growth can be located. Basically, we have $\bm{g_{J}}=(g_{J_{1}},g_{J_{2}},\dots,g_{J_{K}})$ to represent the gradient accumulation of total $K$ human joints. As mentioned in Sec. 4.1, the joints with larger gradients accumulated are more likely to require the growth of child joints. This is achieved by sorting $\bm{g_{J}}$ in descending order $\pi$ to get $\bm{g_{J}}^{\text{sorted}}=(g_{J_{\pi(1)}},g_{J_{\pi(2)}},\dots,g_{J_{\pi(K)}})$.
We set a gradient threshold $\epsilon_{\bm{J}}$ to identify the $J\in\bm{J}_{s}$ that requires growth, following
$\bm{J}_{s}=(J_{\pi(1)},J_{\pi(2)},\dots,J_{\pi(N)})$ (9)
[…]
where we can safely assume $N<K$.
For each joint in $\bm{J}_{s}$, we designate it as a parent joint requiring growth and proceed with the initialization of its child joint. In order to explicitly model each child joint and ensure the consistency with the SMPL paradigm for ease of animation, we maintain an extra joint book $B^{e}=(\text{parent},\bm{j}^{e},\bm{r}^{e})$, that includes the indices of parent joints, the extra joint positions $\bm{j}^{e}$ in canonical space and the extra rotations $\bm{r}^{e}$ in its parent joint coordinate. We initialize the joint position to its parent joint’s position and set the rotation to the identity rotation. The entire parent joint localization and child joint initialization process is guided by the gradients, effectively preventing unnecessary overgrowth and ensuring a dense distribution of the extra joints.
4.3 Extra Joint Optimization
The joint positions and rotations in the extra joint book are set as optimizable, which will be stored and later decoded by two shallow MLPs. According to the SMPL paradigm, the canonical joint position is a time-invariant quantity, therefore the joint position optimization network $\Phi_{p}$ is defined as
$d\bm{j}^{e}(i)=\Phi_{p}(\textit{P.E.}(i)),$ (10)
where P.E. is a positional encoding function as is in [38] and $i$ is the joint index in $B^{e}$. Rotations are also dependent on the timestamp $t$ of each frame, thus the rotation optimization network $\Phi_{r}$ is defined as
$\bm{r}^{e}(i,t)=\Phi_{r}(\textit{P.E.}(i),\textit{P.E.}(t)).$ (11)
Now we have the positions of extra joints $\bm{j}^{e}$ (optimized by the offset $d\bm{j}^{e}$) and the rotations $\bm{r}^{e}$. Each extra joint position is defined in the canonical space, representing the intrinsic properties of the extended skeleton, while the extra joint rotation in parent joint coordinate varies in each frame.
Although both the extra joint positions and rotations are stored in the MLPs, their inputs and outputs are explicit features with real physical meaning, which allows for both implicit and explicit editing. For example, we can interpolate over timestamp $t$ based on $\Phi_{r}$ or directly use explicit inputs to take the place of $\Phi_{r}$ during animation. The MLPs here function as decoders, deriving joint-related values from indices and timestamps, thus effectively circumventing the need for explicit storage of joint values employed in SMPL.
4.4 Inference Process and Training Strategy
In this subsection, we explain how our adaptive growth is integrated with the training process of human gaussians.
First, we initialize the canonical gaussians with standard SMPL vertices. At the beginning of training, we set up a number of warm-up iterations during which no joint growth occurs, and the gaussian fitting is performed following the traditional human gaussian methods. This prevents underfitting due to insufficient training, which could further affect the joint localization in Sec. 4.2.
During the warm-up iterations, canonical human gaussians will be first warped to the observation space according to the LBS weight in Eq. 5. To compensate for the inability of the LBS model to represent gaussians rotation and scale, a deformable network $\Phi_{d}$ is employed to correct the rotation and scale during the LBS process. To distinguish it from the joint rotation $\bm{r}$ of the SMPL human pose, we denote the rotation of the gaussian with subscript $g$, and
$d\bm{r}_{g},d\bm{s}_{g}=\Phi_{d}(\bm{x}_{g},\bm{r}^{0}).$ (12)
The gaussians rotation $\bm{r}_{g}$ and scale $\bm{s}_{g}$ are modified with offsets $d\bm{r}_{g}$ and $d\bm{s}_{g}$ to get the final gaussians in the observation space. In the observation space, we obtain the rendered images through the rasterisation of gaussians and compute the image loss to supervise canonical gaussians.
Once the warm-up iterations completes, we begin the adaptive joint growth. With the MK calculated during the warm-up phase, we can locate the parent joints $\bm{J}_{s}$ that require growth. Then we add grown joints to the extra joint book $B^{e}$, optimizing its positions $\bm{j}^{e}$ and rotations $\bm{r}^{e}$ during the subsequent training. Notably, $\Phi_{\text{lbs}}$ needs to extend its output dimensions to include the blending weights for extra joints, following $K=K^{0}+K^{e}$. Since the extra joints do not have prior LBS weights, their blending weights are entirely learned through $\Phi_{\text{lbs}}$.
Both the warm-up stage and the post-growth learning stage adopt adaptive density control to manage the updates of the gaussians. We also dynamically adjust the threshold of gradient for densification in [25] based on the number of gaussians, in order to balance memory consumption. Please check our supplemental material for details of this design.
5 Experiments
5.1 Dataset
Figure 4: Qualitative comparison on the DNA-Rendering dataset [7] with animatable baselines. We show two cases of hand-held objects (0800_07) and loose-fitting clothing (0811_06) (from top to bottom). Im4d* [33] achieves high-quality rendering but cannot be animated, and we compare it in the supplemental materials. Please check the supplemental video for better visualization.
Our method focuses on hand-held objects and loose-fitting clothing, so we select datasets for experiments accordingly. We notice that the DNA-Rendering dataset [7], by capturing complex scenes of the human body, meets our requirements. Specifically, we selected 8 cases that align with our hypothesis, namely 0041_10, 0090_06, 0176_07, 0800_07 (hand-held objects) and 0007_04, 0014 06, 0051_09, 0811_06 (loose-fitting clothing). For each case, we use 24 surrounding views for training and 6 novel surrounding views for testing. All views are synchronized and contained 100 frames each.
In addition to tackle complicated scenarios, it is essential to ensure the model’s performance in typical scenarios involving tight-fitting clothes. Therefore, we additionally test our method on the ZJU-MoCap [45] dataset. Although the tight-clothing cases are too simple to require joint extension, our overall framework still achieves optimal results. Since this part of the experiment is not directly related to adaptive growth, we refer the readers to the supplemental material for further visualizations.
5.2 Baselines and Metrics
We select the most cutting-edge and representative works from each focus area for a fair comparison. Since 3DGS is currently the leading representation for novel view synthesis, we compare 3DGS-based methods, including 3DGS-Avatar [46], GART [29], and GauHuman [17]. Among them, GART is expected to offer extra animatability for its modeling of implicit global auxiliary bones. Additionally, there is another category of human modeling without incorporating SMPL-like pose priors. Although these methods don’t guarantee an animatable human avatar, they can achieve high-quality rendering, among which, we compare the rendering quality of Im4D [33] with our method.
We conduct a comprehensive comparison of our ToMiE against these methods. We report three key metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [51], and learned perceptual image patch similarity (LPIPS) [56]. Per-scene results can be found in the supplemental material. In addition to comparing the rendering results, we also demonstrate the animatability of the complicated human parts enabled by our method. We strongly recommend readers to watch the supplemental video for a more intuitive understanding of the animating.
5.3 Novel View Synthesis Results
Fig. 4Tab. 1 and Tab. 2 present the results of our method compared to other baselines. In the tables, we showcase two evaluation protocols. The first evaluates the entire image, reflecting the overall rendering quality. The second uses a binary mask to specifically compare the complicated regions, demonstrating how our method outperforms others in these challenging cases. The mask is shown in Fig. 4, whose details can be found in the supplemental material.
Dataset Animatable DNA-Rendering
[…]
3DGS-Avatar        28.93 0.953 0.0459 17.69
GART        29.25 0.958 0.0480 17.12
Im4D        26.28 0.965 0.0308 14.90
GauHuman        30.22 0.962 0.0405 18.26
ToMiE(ours)        31.28 0.966 0.0374 19.75
Table 1: Quantitative comparison between our method and other methods on complicated DNA-Rendering dataset [7]. $\mathbb{B}$ and $\mathbb{G}$ stand for human body and complex clothing (including hand-held objects). We color each result as best, second best and third best. ToMiE achieves the best performance in PSNR and SSIM, while ranking second only to rendering-based Im4D [33] in LPIPS, mainly because rendering-based methods do not consider human structural constraints, resulting in higher visual fidelity.
Dataset Animatable ZJU-Mocap
[…]
3DGS-Avatar        30.61 0.970 0.0296
GART        30.91 0.962 0.0318
Im4D        28.99 0.973 0.0620
GauHuman        30.82 0.962 0.0326
ToMiE(ours)        31.10 0.963 0.0312
Table 2: Quantitative comparison between our method and other methods on ZJU-Mocap dataset [45] with tight-fitting clothing. $\mathbb{B}$ and $\mathbb{G}$ stand for human body and clothing (including hand-held objects). We color each result as best, second best and third best. This comparison demonstrates that ToMiE can achieve comparable (SSIM and LPIPS) or even superior (PSNR) performance to other approaches in tight-fitting scenarios where joint growth is not required.
5.4 Ablation Studies
A. Adaptive Growth Ablation. We remove the adaptive growth to ablate its impact on the rendering results. Tab. 3 “w/o Adaptive Growth” shows a decline in rendering quality, while the hand-held objects and loose-fitting clothing also become not animatable.
B. Non-rigid Design Ablation. In Sec. 4.4, we apply a non-rigid deformation network $\Phi_{d}$ to correct gaussian rotation and scale. As shown in Tab. 3 “w/o Non-rigid Design”, this significantly improves the final rendering quality.
[…]
w/o Adaptive Growth 30.99 0.964 0.0390 19.21
w/o Non-rigid Design 30.78 0.964 0.0363 19.38
Full 31.28 0.966 0.0374 19.75
Table 3: Ablation studies on DNA-Rendering dataset [7]. We independently ablate the adaptive growth strategy and non-rigid design to validate their impact on the overall performance.
5.5 Animating Hand-held Objects and Loose-fitting clothing
We demonstrate the uniqueness of our method, specifically its ability to explicitly animate hand-held objects and loose-fitting clothing. Our animating approach can be implemented in two ways. On the one hand, we can utilize the transformation already recorded in the extra joint book to replay clothing motions. The visualization of this part is shown in Fig. 5. On the other hand, ToMiE also supports bypassing the decoding process of the extra joint book by directly inputting the transformation explicitly. This allows us to customize the motion trajectory of external joints. Due to the limitations of the image’s expressive capabilities, we strongly recommend the readers to watch the supplemental video to check the animated results.
In Fig. 5, we edit the extra joints while keeping the body poses under the SMPL paradigm stationary. Since the implicit auxiliary bones of GART [29] is controlled by the traditional SMPL poses, only the identical appearance can be output when SMPL poses are stationary. In contrast, our method explicitly models hand-held objects and loose-fitting clothing, fully decoupling them from the traditional SMPL poses, enabling free animating.
Figure 5: Animating Results of ToMiE. Our explicit modeling fully decouples hand-held objects and loose-fitting clothing from the human body, enabling part-specific animating.
6 Limitations and Conclusion
Limitations. Although our method enhances the modeling for rigid and non-rigid clothing, it cannot address scenarios involving drastic changes in the topology (e.g., taking off clothes or opening a book). This is because topological changes disrupt the one-to-one correspondence between frames, making the human modeling centered on the canonical space become downgraded. We notice that [41] addresses topology issues by introducing high-dimensional mappings, which could be adapted to build our non-rigid deformation. However, this is not the main scope of this paper and can be explored as future work.
Conclusion. In this paper, we introduce ToMiE, an adaptive growth method designed to extend traditional SMPL skeleton for better modeling of hand-held objects and loose-fitting clothing. In the first stage, we assign the gradient of gaussian points to different joints by combining LBS weights with the motion kernel based on motion priors. This allows us to accurately locate the parent joints that need to grow, avoiding redundant growth. In the second stage, we design an extra joint book to achieve explicit joint modeling and optimize the transformation of the newly grown joints in an end-to-end manner. With the improved designs mentioned above, our ToMiE stands out among numerous state-of-the-art methods, achieving the best rendering quality and animatability of hand-held objects and loose-fitting clothing. We hope the adaptive growing method will spark a renewed discussion on the current capabilities of digital human modeling. What’s more, it is expected to offer some insights for subsequent works related to topology and skeleton generation.
Appendix A Per-scene Results on DNA-Rendering Dataset
We exhibit our per-scene results on DNA-Rendering dataset [7] in Tab. 5. To more clearly demonstrate the results of the growth, we further present the parent joints $J\in\bm{J}_{s}$ for each case in Tab. 4, as defined in Sec. 4.2 (main text). We present more visualizations on the DNA-Rendering dataset in Fig. 8.
Figure 6: Joints distribution we use in our method. We use the SMPL-X model [43] while removing the MANO [47] joints in hands, as we experimentally find that the MANO joints in DNA-Rendering data are inaccurately labeled. More empirically, there is no need to grow extra joints for the fingers.
Sequence Description grown parent joints $\bm{J}_{s}$
0007_04 Waving sleeves [18, 19, 20, 10, 21, 7, 8, 4, 11, 5, 16, 1]
0014_06 Waving sleeves [10, 20, 18, 16, 21, 7, 4, 19, 1, 5, 8, 17, 11, 13]
0041_10 Swinging a feather duster [21, 13]
0051_09 Flowing a dress [ 8, 4, 12, 5, 1, 7, 11, 10, 2]
0090_06 Tending a bonsai [18, 20, 16, 13]
0176_07 Using a hairdryer [21]
0800_07 Watering with a kettle [19, 21, 13, 10]
0811_06 Spinning a dress [4, 7, 2, 1, 5, 0]
Table 4: Description of human action and index of grown parent joints $\bm{J}_{s}$ for each sequence. Please refer to the joint positions in Fig. 6 for a better understanding of the grown joints.
0007_04 0014_06 0041_10
[…]
3DGS-Avatar 28.10 0.949 0.054 18.08 25.40 0.926 0.085 16.09 29.98 0.942 0.054 18.98
GART 28.23 0.952 0.057 17.80 26.87 0.942 0.075 17.16 28.45 0.942 0.060 15.26
Im4D 25.61 0.953 0.037 14.86 26.36 0.955 0.042 16.41 25.75 0.946 0.046 16.51
GauHuman 28.03 0.948 0.054 17.85 27.45 0.941 0.071 17.84 29.41 0.946 0.052 16.14
ToMiE(ours) 29.12 0.953 0.050 18.88 28.72 0.947 0.064 19.46 30.43 0.950 0.048 19.69
0051_09 0090_06 0176_07
[…]
3DGS-Avatar 25.06 0.953 0.041 14.26 32.88 0.968 0.031 19.60 27.93 0.954 0.034 15.99
GART 26.45 0.963 0.042 14.92 34.50 0.975 0.030 20.43 29.87 0.963 0.032 16.38
Im4D 24.36 0.970 0.030 11.99 28.20 0.975 0.023 15.05 26.03 0.982 0.014 13.87
GauHuman 26.24 0.961 0.039 14.75 36.42 0.983 0.021 22.91 32.25 0.980 0.018 19.86
ToMiE(ours) 27.43 0.965 0.037 15.63 36.70 0.984 0.020 22.89 32.77 0.982 0.017 20.29
0800_07 0811_06 average
[…]
3DGS-Avatar 33.57 0.975 0.025 19.64 28.60 0.959 0.044 18.89 28.93 0.953 0.0459 17.69
GART 31.94 0.968 0.035 17.48 27.68 0.961 0.052 17.52 29.25 0.958 0.048 17.12
Im4D 27.56 0.970 0.023 14.88 26.36 0.966 0.032 15.60 26.28 0.965 0.031 14.90
GauHuman 33.48 0.976 0.024 18.87 28.51 0.961 0.046 17.86 30.22 0.962 0.041 18.26
ToMiE(ours) 34.44 0.977 0.021 21.01 30.62 0.967 0.042 20.17 31.28 0.966 0.037 19.75
Table 5: Per-scene quantitative comparisons on the DNA-Rendering [7] dataset.
Appendix B Visualization on ZJU-Mocap Dataset
In Sec. 5.1 (main text), we quantitatively experiment on ZJU-Mocap [45] dataset to validate that our method is also effective in scenarios with tight-fitting garments. In Fig. 7 , we present more visualization results.
Figure 7: Qualitative results on ZJU-Mocap [45] dataset. Please check the zoom-in areas to find that our method reconstructs more details compared to GauHuman [17], even in tight-fitting cases where growth of extra joints is not required.
Figure 8: More qualitative results on the DNA-Rendering [7] dataset. We show cases of hand-held objects (0041_10) and loose-fitting garments (0007_04) (from top to bottom).
Figure 9: Qualitative Comparison of Im4D [33].
Appendix C LBS Visualization
We visualize the LBS weights in Fig. 10, including regions corresponding to maximum and second largest skinning weights.
Figure 10: Visualization of LBS weights.
Appendix D Ablations on Motion Kernel
Our motion kernel aids in the gradient assignment of complex human bodies, helping to accurately identify the joints that need to be expanded. In Fig.3 (main text), the white mask indicates gaussian gradients that belong to the wrist (red point). With MK only, gradients between the wrist and the elbow was incorrectly counted (“misdirect”). With LBS only, the feather duster was partly lost (“misclassify”). So in Eq.(8) (main text), we integrate both to determine parent joints needing growth. We show ablations in Tab. 6 and Fig. 11.
Figure 11: Ablation on Motion Kernel and LBS.
[…]
ToMiE(MK only) 30.27 0.958 0.0432 18.33
ToMiE(LBS only) 30.41 0.960 0.0413 18.52
ToMiE(full) 31.28 0.966 0.0374 19.75
Table 6: Ablation on Motion Kernel and LBS.
Appendix E Training and Rendering Efficiency
ToMiE Gauhuman GART 3DGS-Avatar Im4D
Training Time 30min 25min 15min 20min 3h
Rendering Speed (FPS) 60+ 180+ 60+ 50+ 0.5
Table 7: Average training (till convergence) time and rendering speed on DNA-Rendering dataset [7].
Tab. 7 reports our average training time and rendering speed on DNA-Rendering dataset [7]. We train all these methods on a single GeForce RTX3090.
Appendix F Calculation of Masks with Hand-held Objects and Loose-fitting Garments
Our method focuses primarily on modeling hand-held objects and loose-fitting Garments. In Tab. 1 (main text) and Tab. 2 (main text), we further evaluate the model’s performance in these areas using a binary mask. To generate a mask for each scene that distinguishes regions containing hand-held objects and loose-fitting Garments, we first segment the potential complex garment points in 3D space. Specifically, points predominantly controlled by the extra joints and their parent joints are identified as part of the hand-held objects and loose-fitting Garments. With the pre-trained blending weight, we can easily locate these points, which are then assigned a white color, while all others are marked in black, forming a 3D binary mask. Finally, we obtain a 2D binary mask representing hand-held objects and loose-fitting Garments by applying Gaussian rasterization to the 3D mask. Since we rely solely on this binary mask for metrics evaluation, this post-processing method for calculating masks is permissible.
Appendix G Adjustment of the Gradient Threshold for Densification
In scenarios with hand-held objects and loose-fitting garments, to prevent excessive gaussian points from causing high memory consumption, we propose an adaptive suppression strategy to keep the number of gaussian points within a reasonable range. This is achieved by dynamically adjusting the threshold of the gradient for densification $\epsilon_{d}$ in [25]. This threshold $\epsilon_{d}$ represents that points with accumulated gradients exceeding it will be densified. Therefore, a larger threshold results in fewer split points, and vice versa.
Let us assume the desired maximum number of gaussian points is $N$. After each iteration, if the current number of gaussian points $n$ exceeds $N$, we will increase $\epsilon_{d}$ according to
$\epsilon_{d}=(a+\frac{n-N}{b})\epsilon_{d_{0}}.$ (13)
In the practical implementation, we set $N=3\times 10^{4}$, $a=2$, $b=5\times 10^{3}$, and $\epsilon_{d_{0}}=5\times 10^{-4}$.
Appendix H Details of Hyperparameters
The $\lambda$ of balancing MK weight and LBS weight in Eq. (8) (main text) is set to $0.4$. The gradient threshold $\epsilon_{\bm{J}}$ to identify the $J\in\bm{J}_{s}$ that requires growth in Eq. 13 is set to $3.5\times 10^{-6}$. For network hyperparameters, we detail the number of layers and the width of the MLP network design. $\Phi_{\text{lbs}}$ has $D=4$ and $W=128$. $\Phi_{p}$ has $D=4$ and $W=256$. $\Phi_{r}$ has $D=4$ and $W=128$. $\Phi_{d}$ has $D=2$ and $W=128$. Specifically, we initialize the weights of the last layer in $\Phi_{p}$ and $\Phi_{r}$ as a tiny value $1\times 10^{-2}$. This provides stability for the initial training phase. Warm-up iterations number in Sec. 4.4 (main text) is set to $8\times 10^{3}$.
Appendix I Visual Comparison with Im4D
As Im4D [33] can only perform rendering, we didn’t compare its visual quality with other animatable methods in the main text. In Fig. 9, we compare its rendering quality with our ToMiE. Although it is purely rendering-based and not constrained by SMPL skeleton, ensuring high visual fidelity (LPIPS), it suffers from the inability to animate and color distortion (PSNR).
Appendix J Supplemental Video
Our supplemental video consists of three parts. First, we present monocular rendering results to demonstrate that we can accurately render hand-held objects and loose-fitting garments on the human body. Next, we perform 360-degree rendering of the full video to validate our generalization capability on novel views. These two parts are visually compared with GauHuman [17], one of the current state-of-the-art methods. Finally, we fix standard human skeleton still, animating extra joints to show our decoupling and explicit animating capability. Since image quality may not fully demonstrate the effectiveness of human reconstruction, especially for animating results, we recommend that readers refer to this supplemental video for better visualization.